OpenStudy (anonymous):

4x-y=4 x+4y=18 a) 2 4 b) 4 2 c) 1 4 d) 4 1

OpenStudy (saifoo.khan):

x = 2 y = 4

OpenStudy (anonymous):

option a

OpenStudy (anonymous):

what about this ? niti Group Star x+3y=-7 2x+4y=-4 a) 3y-7 b)-3y-7 c)3y+7 d)-3y+7

OpenStudy (saifoo.khan):

x = 8 y = -5

OpenStudy (anonymous):

y=? 4x-y=4 x+4y=18 /(x4) ------------------- -17y=-68 y=4 x=? 4x-y=4 /(x4) x+4y=18 ------------------ 17x=34 x=2 Option A